# Searching for gravitational waves from stellar-mass binary black holes early inspiral

### v1

@@ V.1 Detection
Figure 9 showcases the detection ability of the CNN model, which has been trained with SNR uniformly distributed within the 30 to 50 range, to differentiate signals based on their varying SNR. It is observed that an increase in the SNR corresponds to an augmentation in the TAP, with the specifics delineated in Table 4. The final trained detection CNN model can effectively detect sBBH sources with an SNR of 50 or higher, achieving an area under the curve (AUC) of 0.961. Additionally, it demonstrates that sources with fixed SNRs of 40, 30, and 20 achieved corresponding area under the curves values of 0.847, 0.669, and 0.557, respectively.
@@ V.2 Mass estimation
In order to estimate the value of the chirp mass by the point parameter estimation CNN model trained with a uniformly sampled SNR ranging from 30 to 50, we took 2000 test samples with a fixed SNR of 50 which had the same parameter distribution as the training samples. As shown in Figure 10, most of the chirp mass predictions are tightly distributed along the diagonal, indicating that there is consistency between the predicted values and actual values of the source’s chirp mass. In this test dataset, 90% of the samples have an absolute error of 2.48 $M_{\odot}$ and a relative error of 0.12. This means that our model can predict the chirp mass of SNR 50 signals in TianQin runs. Leaving aside those sources with larger errors, the significant variance observed in most sources yielding accurate predictions can largely be ascribed to added detector noise.
We also studied the characteristics of signals with larger prediction errors for chirp mass (absolute error exceeding 10 solar masses). We found that some signals have a smaller fitting factor, indicating that more information was lost after the IPCA projection. Signals characterized by a larger symmetric mass ratio or larger effective spins will exhibit a reduced number of data points in frequency domain within the same duration. This particular property introduces challenges for CNNs as they strive to effectively learn these signal features and make accurate predictions of the chirp mass. From the current results, the point estimation CNN can still quickly provide chirp mass predictions for most sBBH signals of SNR 50 with a good accuracy.
@@ VI Conclusion and discussion
In this study, we effectively compressed the data through low rank matrix approximation, and extracted the main features of the frequency signal amplitude, thereby building the IPCA model. By projecting the simulated detector data (including signal plus noise samples and pure noise samples) through IPCA, we classify the compressed data into two categories by a trained CNN and show the sensitivity via the ROC curve. Simultaneously, in terms of the point estimation of parameters, we utilized the CNN model to estimate chirp mass, marginalising over all other parameters, for the detector data containing sBBH signals with a SNR of 50, and estimated its measurement error (90% confidence interval) to be 2.48 solar masses.
@@
In conclusion, this strategy of combining data compression and search presents a promising new approach for the detection of long-lived sBBH signals. Intending to refine this method further in the future, our search for early inspiral phases of sBBH will directly contribute to a better understanding of the formation channels of sBBH.

### v2

@@ V.1 Detection
Figure 9 showcases the detection ability of the CNN model, which has been trained with SNR uniformly distributed within the 30 to 50 range, to differentiate signals based on their varying SNR. It is observed that an increase in the SNR corresponds to an augmentation in the TAP, with the specifics delineated in Table 4. The final trained detection CNN model can effectively detect sBBH sources with an SNR of 50 or higher, achieving an area under the curve (AUC) of 0.958. Additionally, it demonstrates that sources with fixed SNRs of 40, 30, and 20 achieved corresponding area under the curves values of 0.847, 0.671, and 0.558, respectively.
@@ V.2 Mass estimation
In order to estimate the value of the chirp mass by the point parameter estimation CNN model trained with a uniformly sampled SNR ranging from 30 to 50, we took 2000 test samples with a fixed SNR of 50 which had the same parameter distribution as the training samples. As shown in Figure 10, most of the chirp mass predictions are tightly distributed along the diagonal, indicating that there is consistency between the predicted values and actual values of the source’s chirp mass. In this test dataset, 90% of the samples have an absolute error of 2.49 $M_{\odot}$ and a relative error of 0.13. This means that our model can predict the chirp mass of SNR 50 signals in TianQin runs. Leaving aside those sources with larger errors, the significant variance observed in most sources yielding accurate predictions can largely be ascribed to added detector noise.
We also studied the characteristics of signals with larger prediction errors for chirp mass (absolute error exceeding 10 solar masses). We found that some signals have a smaller fitting factor, indicating that more information was lost after the IPCA projection. Signals characterized by a larger symmetric mass ratio will exhibit a reduced number of data points in frequency domain within the same duration. This particular property may introduce challenges for CNNs as they strive to effectively learn these signal features and make accurate predictions of the chirp mass. From the current results, the point estimation CNN can still quickly provide chirp mass predictions for most sBBH signals of SNR 50 with a good accuracy.
@@ VI Conclusion and discussion
In this study, we effectively compressed the data through low rank matrix approximation, and extracted the main features of the frequency signal amplitude, thereby building the IPCA model. By projecting the simulated detector data (including signal plus noise samples and pure noise samples) through IPCA, we classify the compressed data into two categories by a trained CNN and show the sensitivity via the ROC curve. Simultaneously, in terms of the point estimation of parameters, we utilized the CNN model to estimate chirp mass, marginalising over all other parameters, for the detector data containing sBBH signals with a SNR of 50, and estimated its measurement error (90% confidence interval) to be 2.49 solar masses.
@@
From a broad perspective, the sensitivity of our pipeline is more efficient than the excess power method, which only uses amplitude. This is because we incorporate additional signal information during compression. The extracted bases encapsulate the correlation of data points in signals - if the detector records data in the presence of a signal, this correlation becomes evident during compression. This compression technique could also enhance our understanding of the latent space for sBBH signals or it could be employed to assess the reduced bias across the entire template bank, as demonstrated in the paper [62]. Moreover, this strategy of combining data compression and search presents a promising new approach for the detection of long-lived sBBH signals. Intending to refine this method further in the future, our search for early inspiral phases of sBBH will directly contribute to a better understanding of the formation channels of sBBH.
@@ Acknowledgements.
Appendix A IPCA
We extracted 480 basis for the IPCA models, constrained by the processing capacity of our computational cluster, notably the GPU memory. This appendix was done to demonstrate the influence of the number of training signals when constructing a compression model using IPCA. In the figure 11, the variance ratio for different IPCA models within the A-channel signal exhibits minimal change. Concurrently, the basis also registered minor adjustments.
Figure 11: Variance ratio plotted against the number of training signals for the IPCAA model.